OpenStudy (anonymous):

Given the points (2, 3) and (6, 8), write the point-slope form of the equation. y - 3 = x - 2 y - 3 = 5/4 (x - 2) y - 8 = 4/5 (x - 6)

hartnn (hartnn):

first find slope given 2 points

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

@hartnn so is it y - 8 = 4/5 (x - 6)?

hartnn (hartnn):

what did u get your slope as ?

OpenStudy (anonymous):

@hartnn -2?

hartnn (hartnn):

how ?

OpenStudy (anonymous):

@hartnn by making a graph and putting the points on it then getting the line on it